RC Meeting and Run Plan, Tuesday, February 26

From HallCWiki
Jump to: navigation, search

Main | Online | RC | Run Plan | Official Website | Redmine | Index | Nav
Jpsi007.png

RC Meeting Information

RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Good beam overnight. Finished with ~2 hrs on Dummy target for Phase-1 kinematics.
  • "RF Recovery" started at 8. Went to controlled access and Steve did some work in the hall.
  • Rotated and set magnets for Phase-2 kinematic
  • Beam came back around 1:30pm and pretty bad since.
  • Will continue running J/Psi on kinematic Phase-2.

Current Run Plan

  • J/Psi Experiment:
  • Run Production Phase-2. Shooting for a total of 6.5 C for phase-2 kinematic before returning to Phase-3.

Notices in Effect for Shift Crew

  • Shift Leader is responsible for 50k replays. Even if the TO is actually doing them.
  • Target Operators: When the alarm handler goes white, it is disconnected from the IOC. Log it and include carlin@jlab.org every time. Acknowledge the warning. If it is still white and not coming back, restart the alarm handler. Do not permanently silence the alarm handler.
  • Missing Reference Times during the replay: check in the 50k. Check in full replay, if able. If it happens, call RC. Guidance to solve it is here: https://logbooks.jlab.org/entry/3654584


Current Issues

From Steve's logbook entry:

  • Concerning target alarms (loss of cooling power/over head) continued when running with hall A.

Planned Accesses to the Hall

Planned Accelerator Configuration Changes

Special Requests

  • Latif's GEM testing completed this morning during beam studies.